sandra.is starting a vacuum cleaning business. her start up cost to buy a vacuum and other equipment is $550 she determined she will use about $8 in cleaning supplies for each house. if she charges $58 per house what is her break even point
Answer by ikleyn(52858)

.
sandra.is starting a vacuum cleaning business. her start up cost to buy a vacuum and other equipment
is $550 she determined she will use about $8 in cleaning supplies for each house.
if she charges $58 per house what is her break even point
~~~~~~~~~~~~~~~~


    550 + 8x = 58x

    
    550 = 58x - 8x

    550 = 50x

     x  = 550/50 = 11.    ANSWER

Solved.

The formulas and the starting equation are self-explanatory.